OpenStudy (anonymous):

Let f(x)= 3x^2 / (2x+1) . At what value of x does f(x) have a tangent with slope 4/3?

OpenStudy (anonymous):

did you find the derivative?

OpenStudy (anonymous):

i found the derivative to be \[6x ^{2} + 6x \div (2x+1)^{2 }\]

OpenStudy (anonymous):

you should get \[\frac{6x^2+6x}{(2x+1)^2}\] so solve \[\frac{6x^2+6x}{(2x+1)^2}=\frac{4}{3}\] via \[3(2x+1)^2=3(6x^2+6x)\]

OpenStudy (anonymous):

ok typo, i meant \[4(2x+1)^2=3(6x^2+6x)\]

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

yw annoying quadratic, but you can solve it yes?